You are fascinated by the genetics of butterflies, and are especially intrigued by the fact that in butterflies, females are the heterogametic sex. you cross true-breeding yellow female butterflies with eyespots on their wings to true-breeding red males without eyespots. You know that the F1 progeny were all orange, but your assistant left the notes about presence or absence of eyespots out in the rain, and the data were lost. you decide to self the F1 to confirm your suspicion about the inheritance of color, and to see if you can figure out what is going on with the inheritance of eyespots. Your F2 progeny data are presented in the table below.

Yellow, spotted Red, spotted orange, spotted yellow, no spots Red, no spots Orange, No spots
F2 Males 16 17 31 47 46 94
F2 Females 17 16 31 46 47 94

1. How is wing color inherited?

A. Red completely dominant to yellow.

B. Yellow codominant with red.

C. Yellow completely dominant to red.

D. Red incompletely dominant to yellow.

2. What proportion of the F2 male butterflies are yellow?

A. 1/2

B. 3/4

C. 1/4

D. 2/3

E. 1/3

3. How is the trait of spots on things inherited?

A. No spots codominant with spots.

B. Spots incompletely dominant to no spots.

C. No spots completely dominant to spots.

D. Spots completely dominant to no spots.
Answer #1

As there is mixing of color Red and Yellow orange so it is incomplete dominance example and by seeing the ratio in orange, No

1.D. Red is incomplete dominant over yellow.

as in incomplete dominance allelic mixing produre mixed character like red mix with yellow to give orange.

2.C.1/4

F2 male yellow are present in yellow spotted(1) and yellow no spots(3)

soproportion will be =( 1 + 3 ) / (1+1+2+3+3+6) = 4/16 =1/4

3.C.no spos completely dominant to spots